OpenStudy (rock_mit182):

medal for this one! if on the expression of X2Y, the value of x is diminished in a 40% and the value of Y in a 25%, then the value of the expression would be diminished in a ?

OpenStudy (rock_mit182):

the expression is\[x ^{2}y\]

OpenStudy (mathmale):

Your expression is (x)^2*(y). Why not replace this (x) with (0.60x) and this (y) with (0.75y) and see waht you get through simplification? Why 0.60 instead of 0.40, and why 0.75 8nstead of 0.25?

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

thx for showing the work

OpenStudy (rock_mit182):

so the (new x) = x-0.4x = 0.6x the (new y) = y-0.25y = 0.75y now i think i have to plug and chuff in the expression . .

OpenStudy (mathmale):

Right on. See what you can do. Please share your results.

OpenStudy (rock_mit182):

would be like that ? \[x ^{2}y-(0.6x)^{2}0.75y\]

OpenStudy (rock_mit182):

it doesn't make sense to me the options are different

OpenStudy (rock_mit182):

45% 32,5% 27% 22,5%

OpenStudy (rock_mit182):

@mathmale

OpenStudy (mathmale):

You are starting with x^2*y, right? If you diminish x by 40%, you'll then have 0.6 x. If you dim. y by 25%, you'll then have 0.75 y. Now evaluate (x_^2*y, using these new values 0.6x and 0.75y. Factor x^2*y out of the result. What's left is a coefficient. Figure out the value / size of that coeff., and then you'll have your answer. Good luck!

OpenStudy (rock_mit182):